A triangle has sides with lengths of 5 feet, 12 feet, and 13 feet. Is it a right triangle? yes or no

Found 3 solutions by stanbon, Amrata1, MathTherapy:
Answer by stanbon(75887)

A triangle has sides with lengths of 5 feet, 12 feet, and 13 feet. Is it a right triangle? yes or no
----
Ans:: yes if 13^2 = 12^2 + 5^2
yes if 169 = 144 + 25
yes if 169 = 169
---------------
Ans: Yes
------------
Cheers,
Stan H.
-------------

Answer by Amrata1(11)

It is a right angled triangle.
You need to check whether Pythagoras theorem can be applied or not.
If suppose it is a right angled triangle,
Then hypotenuse will be the longest side.
Here it is 13 ft.
So check whether 5^2 + 12^2 = 13^2 .
Since it is true, hence triangle is a right angled triangle.

Answer by MathTherapy(10556)


A triangle has sides with lengths of 5 feet, 12 feet, and 13 feet. Is it a right triangle? yes or no
A triangle with 5-12-13 sides is considered a Pythagorean triple, and represents one of many right-angled triangles.